The following is total monthly budgeted cost and activity information for the four activity centers in the billing department of Oregon Power Company:

Activity Center

Variable

Fixed

Cost Driver

Account inquiry

$79,584

$155,000

3,200 labor hours

Correspondence

$9,947

$22,000

2,900 letters

Account billing

$175,000

$71,000

2,500,000 lines

Bill verification

$9,020

$80,000

20,500 accounts


In September, actual costs and activity were as follows:

Activity Center

Total Costs

Driver Amount

Account inquiry

$240,167

3,400 labor hours

Correspondence

$31,062

2,820 letters

Account billing

$236,262

2,360,000 lines

Bill verification

$88,456

20,340 accounts


Required
Compute the flexible-budget variances for the following two activity cost items (round all answers to the nearest dollar and enter favorable variances as positive numbers and unfavorable variances as negative numbers):

  Correspondence    

  Bill verificatio
Answer #1
Flexible-budget
Correspondence     31673 =22000+(9947/2900*2820)
Bill verification 88950 =80000+(9020/20500*20340)
Flexible-budget variances
Correspondence     611 =31673-31062
Bill verification 494 =88950-88456